A motorist travelling with a constant speed 15m/s passes a school crossing corner where speed limit is 10m/s just as the motorist passes a police on a motorcycle starts off in pursuit with a constant acceleration 3m/s2 .How much time elapses before he catches up with the motorist?
What is the officers speed? What is the total distance travelled byveach vehicle at that point?
1
Expert's answer
2020-02-27T10:10:32-0500
"d=vt=0.5at^2\to v=0.5at"

"15=0.5(3)t"

"t=10\ s"

"u=at=3(10)=30\frac{m}{s}"

"d=vt=(15)(10)=150\ m"